# Penalty decomposition derivative free method for the minimization of partially separable functions over a convex feasible set

### v1

@@ 5 Computational experiments
The termination criterion for PDDF consists in the conditions (22) with $\xi_{k}=10^{-4}$ for all $k$, whereas for LS we employ $\max_{i=1,\dots,n}\alpha_{i}\leq 10^{-4}$. In case one algorithm does not meet the termination condition within 2 hours of wall-clock time or 10,000 outer cycle iterations, we stop the experiment. The initial stepsize is set to $1$ for both methods. As regards the line search, both methods use $\gamma=10^{-6}$ and $\theta=0.5$. For PDDF we set $\tau_{0}=f(x^{0})/(10^{2}m)$ and $\tau_{k+1}=\min\{1.05\tau_{k},f(x^{0})/m\}$, with $x^{0}\in\mathbb{R}^{n}$ being the initial solution of the problem considered. The upper bound on the penalty parameter is set for numerical stability reasons. The computational experiments here reported were implemented in Python3 and executed on a computer running Ubuntu 22.04 with an Intel Xeon Processor E5-2430 v2 6 cores 2.50 GHz and 32 GB RAM. The implementation of the proposed PDDF method is available at github.com/dadoPuccio/PD_derivative_free.
In order to offer a fair comparison between the two approaches, we count the number of function evaluations of each function $f_{j}$ that composes the objective function. By assuming that the computational burden for evaluating each $f_{j}$ is similar, the total of individual sub-function evaluations provides an unbiased metric to compare PDDF and LS. Table 2 reports the total count for PDDF and LS when their respective termination conditions are met. We observe that in most cases PDDF outperforms LS as it converges using fewer function evaluations. We also notice that LS scales very poorly with problem dimensionality: exploiting the problem structure, as in PDDF, does become crucial in this setting. Problem DIXMAANI results particularly challenging for LS, that is not capable to trigger the termination condition before the time limit is reached at any of the three largest dimensions we tested. In Table 3 we report the function values obtained by PDDF and LS when the termination condition is met. We observe that both algorithms are capable of obtaining high quality solutions, with the only exception being LS in the largest ROSENBR problem, as in that case the time limit is hit before the algorithm is capable to get a satisfactory solution.
Table 2: Total number of function evaluations for PDDF and LS when the termination condition is met. We use MAX_T and MAX_I to indicate cases where the maximum execution time or the maximum number of iterations was reached.
Problem $\mathbf{n}=n_{1}$ $\mathbf{n}=n_{2}$ $\mathbf{n}=n_{3}$ $\mathbf{n}=n_{4}$ $\mathbf{n}=n_{5}$ $\mathbf{n}=n_{6}$
PDDF LS PDDF LS PDDF LS PDDF LS PDDF LS PDDF LS
ARWHEAD 810 2709 4410 7.4e4 8910 3.0e5 4.5e4 7.5e6 9.0e4 3.0e7 4.5e5 7.5e8
BDEXP 7344 4424 4.2e4 1.4e5 8.6e4 5.6e5 4.4e5 1.4e7 8.7e5 5.7e7 4.4e6 1.4e9
BDQRTIC 4.1e4 3810 3.1e5 1.7e5 6.6e5 7.0e5 3.4e6 2.0e7 6.8e6 8.0e7 3.4e7 2.2e9
BEALES 600 2055 3000 5.1e4 6000 2.1e5 3.0e4 5.1e6 6.0e4 2.1e7 3.0e5 5.1e8
BROYDN3D 2.7e4 7200 1.8e5 1.6e5 4.2e5 6.4e5 2.1e6 1.6e7 4.1e6 6.4e7 2.1e7 1.6e9
DIXMAANA 3375 8640 1.1e4 1.0e5 2.3e4 4.0e5 1.2e5 9.6e6 2.4e5 3.8e7 1.3e6 9.6e8
DIXMAANI 6300 8625 2.1e4 1.0e5 5.2e4 4.0e5 3.0e5 MAX_T 1.7e6 MAX_T 1.3e7 MAX_T
ENGVAL 1.2e4 6174 6.5e4 1.8e5 1.3e5 7.3e5 6.5e5 1.8e7 1.3e6 7.3e7 6.5e6 1.8e9
MOREBV 1872 9.7e4 8372 3.8e5 1.6e4 9.2e5 7.2e4 9.6e6 1.4e5 3.8e7 6.7e5 7.5e8
NZF1 4875 3990 3.7e4 4.2e4 3.7e5 1.6e6 2.3e6 6.4e7 3.9e5 2.5e8 MAX_T MAX_T
POWSING 1160 3605 3016 2.4e4 5800 9.0e4 2.9e4 2.3e6 5.8e4 9.0e6 2.9e5 2.3e8
ROSENBR 4.3e4 7.3e4 2.2e5 1.8e6 4.3e5 7.3e6 2.2e6 1.8e8 4.3e6 7.3e8 2.2e7 MAX_T
TRIDIA 1.7e4 7350 1.3e5 2.5e5 4.3e5 1.1e6 1.1e7 2.7e7 4.7e7 1.1e8 MAX_T 2.8e9
WOODS 3690 2.3e4 7380 9.1e4 3.7e4 2.3e6 7.4e4 9.1e6 3.7e5 2.3e8 7.4e5 9.1e8
Table 3: Objective function values obtained with PDDF and LS when the termination condition is met.
Problem $\mathbf{n}=n_{1}$ $\mathbf{n}=n_{2}$ $\mathbf{n}=n_{3}$ $\mathbf{n}=n_{4}$ $\mathbf{n}=n_{5}$ $\mathbf{n}=n_{6}$
PDDF LS PDDF LS PDDF LS PDDF LS PDDF LS PDDF LS
ARWHEAD 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
BDEXP 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
BDQRTIC 11.9 11.9 106.0 106.0 223.7 223.7 1.2e3 1.2e3 2.3e3 2.3e3 1.2e4 1.2e4
BEALES 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
BROYDN3D 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
DIXMAANA 15.0 15.0 51.0 51.0 102.0 102.0 501.0 501.0 1.0e3 1.0e3 5.0e3 5.0e3
DIXMAANI 15.0 15.0 51.0 51.0 102.0 102.0 501.0 501.0 1.0e3 1.0e3 5.0e3 5.0e3
ENGVAL 9.2 9.2 53.6 53.6 109.1 109.1 553.1 553.1 1.1e3 1.1e3 5.5e3 5.5e3
MOREBV 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
NZF1 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
POWSING 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
ROSENBR 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 74.2
TRIDIA 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.1 0.0
WOODS 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
@@
We provide now the reader with some results concerning the wall-clock time that PDDF method employs to converge. Given the structure of the subproblems obtained with the penalty decomposition scheme, each duplicate vector of variables can be optimized independently of each other: therefore, a parallel implementation of the penalty decomposition derivative free method (PDDFP) can be fruitfully employed. This is particularly appealing for real-world black-box problems, where objective function evaluations are typically expensive to compute; hence, being able of evaluating multiple functions $f_{j}$ in parallel allows for considerable improvements to the overall computational time. The LS method, on the other hand, cannot exploit parallel computation; for a more challenging test, PDDFP is thus compared with a structure-aware implementation of the classical line-search based derivative-free coordinate descent method (SALS): by keeping track of the values of the individual functions $f_{j}$, it is possible to evaluate only a subset of the functions when a new point is considered, i.e., those impacted by changes to the currently considered variable. As regards the parallel implementation, we limit the maximum number of workers to 12. We simulate the high cost of function evaluations by introducing a 10ms pause at each evaluation. In Table 4 we report the outcome of this analysis. We observe that PDDFP consistently outperforms SALS when the problem dimension is large, as in that case the number of functions $f_{j}$ that constitute the objective is larger and the parallel implementation can be fruitfully employed.
Table 4: Wall-clock time required to reach the termination condition for the parallel version of PDDF and the structure-aware LS. Time is expressed in seconds.
Problem $\mathbf{n}=n_{1}$ $\mathbf{n}=n_{2}$ $\mathbf{n}=n_{3}$ $\mathbf{n}=n_{4}$
PDDFP SALS PDDFP SALS PDDFP SALS PDDFP SALS
ARWHEAD 1.1 5.6 3.0 30.4 5.1 61.5 21.8 311.1
BDEXP 11.2 13.7 21.6 83.8 48.2 171.4 204.1 873.6
BDQRTIC 79.7 19.3 159.0 174.7 310.7 352.4 1.5e3 2.0e3
BEALES 1.4 4.3 2.7 21.3 4.1 42.5 15.5 213.5
BROYDN3D 37.4 20.6 103.5 97.6 220.3 195.7 970.0 982.5
DIXMAANA 3.7 11.8 6.4 40.2 11.5 80.4 57.0 395.9
DIXMAANI 7.4 11.8 11.4 40.2 25.5 80.4 143.2 MAX_T
ENGVAL 18.0 12.7 40.9 73.9 72.0 149.3 309.3 754.6
MOREBV 2.0 14.8 5.3 72.3 8.4 138.3 33.7 495.3
NZF1 16.5 12.0 42.7 40.1 238.3 433.0 1.2e3 3.5e3
POWSING 2.6 7.2 2.7 18.8 5.1 36.2 15.8 181.5
ROSENBR 96.1 149.2 189.6 746.2 290.3 1.5e3 1.1e3 MAX_T
TRIDIA 23.0 14.4 75.6 100.1 222.5 215.8 5.1e3 1.1e3
WOODS 2.7 19.7 4.3 39.3 18.3 196.9 36.0 394.4
@@ 6 Conclusions
In this paper, we described and formally analyzed a penalty decomposition scheme for derivative-free optimization of coordinate partially separable finite-sum problems. The approach is proved to be well defined and possess global convergence guarantees under very mild standard assumptions. The results of preliminary computational experiments indicate that the proposed strategy shall be far superior than the blind employment of classical coordinate search approaches, especially as the size of the problems grows. Moreover, the approach is well-suited for parallelization; in scenarios where multi-processing is possible and the cost of computation is dominated by the evaluation of the black-box function, the proposed framework is shown to be superior also to a structure-informed smart implementation of the line-search based coordinate descent method.
The nice feedback obtained by the preliminary experiments opens up to interesting paths that could be explored in future research. First, working independently with each term in the objective should make the entire process less vulnerable to the presence of noise in a subset of the objectives. Moreover, the separability with respect to the sum terms is particularly nice in mixed setting where derivatives are available for some of the objectives [18]; indeed, gradient based optimization steps could be carried out w.r.t. blocks of variables associated with white-box terms; this would not be possible directly tackling the original formulation of the problem, where the same variables possibly appear in white and black box terms simultaneously. Finally, a strength point of penalty decomposition approaches in general is that they are able to handle additional constraints with a limited overhead (see, e.g., the discussion in [13]): indeed, additional constraints can be moved to the penalty function and handled naturally within the sequential penalty scheme.
@@
Matteo Lapucci: $0$-0002-2488-5486
Davide Pucci: $0$-0003-4424-9185
Marco Sciandrone: $0$-0002-5234-3174

### v3

@@ 5 Computational experiments
The inner termination criterion for PDDF consists in the conditions (21) with $\xi_{k}=10^{-2}$ for all $k$. The outer loop is then stopped when either $\|x^{k+1}-x^{k}\|\leq 10^{-2}$ or $k=100$. The LS method is then run starting from the obtained solution for a refinement process. For LS, used either standalone or as a refiner following PDDF we employ $\max_{i=1,\dots,n}\alpha_{i}\leq 10^{-4}$ as stopping condition. Similarly, MADS terminates when the mesh size is below $10^{-4}$. In case one algorithm does not meet the termination condition within 10 minutes of wall-clock time we terminate the experiment. The initial stepsize is set to $1$ for both PDDF and LS, while MADS uses the default initial mesh size. As regards the line-search, both PDDF and LS use $\gamma=10^{-6}$ and $\theta=0.5$. For PDDF we set $\tau_{0}=1$ and $\tau_{k+1}=\min\{1.1\tau_{k},10^{8}\}$. When considering constrained problems, PDDF handles the box explicitly in the line-search procedure, as LS does. The computational experiments here reported were implemented in Python3 and executed on a computer running Ubuntu 24.04 with an Intel Core i5-13400F 10 cores 4.60 GHz processor and 32 GB RAM. The implementation of the proposed PDDF method is available at github.com/dadoPuccio/PD_derivative_free.
In order to offer a fair comparison between the three approaches, we count the number of function evaluations of each function $f_{j}$ that composes the objective. By assuming that the computational burden for evaluating each $f_{j}$ is similar, the total of individual sub-function evaluations provides an unbiased metric to compare PDDF, LS and MADS. As a tool for comparing the behavior of the algorithm considered we employ both data profiles [27] and performance profiles [11]. Data profiles report the percentage of solved problems as the computational budget increases, while, given a metric, performance profiles report the fraction of fraction of problems solved within a given performance ratio with respect to the best solver. We account a problem as solved by the solver $S$ in case $f(x_{0})-f(x_{S})\geq(1-\epsilon)(f(x_{0})-f(x_{\text{best}}))$, where $f(x_{0})$ is the objective at the starting point, $x_{S}$ is the solution found by the solver $S$ and $x_{\text{best}}$ is the best solution found by all optimizers. In data profiles we normalize the number of function evaluations in problems with different dimension $n$ and different number of functions $m$ by considering groups of $m\times(n+1)$ sub-function evaluations. All problems are initialized to the CUTEst default starting solution for all solvers.
Figure 1: Data profiles and performance profiles in the problems derived from CUTEst.
In Figure 1 we report the data profiles and performance profiles in the set of problems derived from CUTEst. We consider unconstrained and bound-constrained problems separately in the profiles. In both cases we can observe that PDDF is often capable of obtaining good quality solutions with fewer function evaluations, especially if the required precision is not extremely high (as often happens in black-box settings). On the contrary MADS is far behind the others, probably due to the large problem dimensionalities considered in our experimental setting.
Figure 2: Performance profile of the wall-clock time require do reach the termination condition in the problems derived from CUTEst.
@@
We provide now the reader with some results concerning the wall-clock time that PDDF method employs to converge. Given the structure of the subproblems obtained with the penalty decomposition scheme, each duplicate vector of variables can be optimized independently of each other: therefore, a parallel implementation of the penalty decomposition derivative free method (PDDFP) can be fruitfully employed. This is particularly appealing for real-world black-box problems, where objective function evaluations are typically expensive to compute; hence, being able of evaluating multiple functions $f_{j}$ in parallel allows for considerable improvements to the overall computational time. The scaling properties of PDDFP are tightly linked to the number of sub-functions that compose an objective function. Given duplication of variables for each sub-function, adding a processor will be beneficial for PDDFP only in case the total number of computing units stays below the number of sub-functions.
The LS method, on the other hand, cannot exploit parallel computation; for a more challenging test, PDDFP is thus compared with a structure-aware implementation of the classical line-search based derivative-free coordinate descent method (SALS): by keeping track of the values of the individual functions $f_{j}$, it is possible to evaluate only a subset of the functions when a new point is considered, i.e., those impacted by changes to the currently considered variable.
As regards the parallel implementation, we limit the maximum number of workers to 12. We simulate the high cost of function evaluations by introducing a 1ms pause at each evaluation. In 2 we report the outcome of this analysis in form of performance profiles. We observe that PDDFP is superior to SALS especially in unconstrained problems. For bound-constrained problems, the gap between the two algorithms is smaller because there are fewer cases with a large number of sub-functions $m$, which limits PDDFP’s ability to fully exploit the available parallel workers.
In the final part for our computational analysis, we compare the behavior of the considered algorithms in two different modification of the Lockwood problem [17]. As earlier discussed in this Section, the original formulation of the problem cannot be handled by PDDF because of the structure of its constraints. We devise and adapted variant of (23) where the constraints are moved to the objective function in the form of static penalties, leading to
$\min_{x}\left\{f(x)=\sum_{j=1}^{6}x_{j}+\max\{c_{1}(x),0\}+\max\{c_{2}(x),0\}:x\in[0,2\cdot 10^{4}]^{6}\right\},$ (L1)
leading to a 6-dimensional problem with box-constraints and an objective function made up of three terms. The resulting problem does fit in our framework but the three sub-function that compose the objective are not partially separable. We therefore consider a partially separable version of the problem (L1) by considering four copies of its objective function with different (random) weights assigned to the three terms, and sharing the last variable of one copy as the first variable of the next one. The resulting problem, that we indicate with (L2), is a 21-dimensional box-constrained problem with a partially separable objective function mad up of four terms.
Figure 3: Evolution of the objective function value with respect to the total number of sub-function evaluations for the compared optimization algorithms in the two versions of the Lockwood problem considered.
In this case the termination condition of PDDF is $\xi_{k}=10^{-1}$ and $\|x^{k+1}-x^{k}\|\leq 10^{-1}$, followed by LS to further refine the solution. For LS we employ $\max_{i=1,\dots,n}\alpha_{i}\leq 10^{-2}$ as the termination criteria, while MADS terminates when the mesh size is below $10^{-2}$. In Figure 3 we report the evolution of the objective function during the optimization when considering PDDF, LS and MADS. In the case of (L1) the penalty decomposition scheme of PDDF is disfavored by the not-partially separable nature of the problem, therefore leading to the worst performing algorithm in this case. On the other hand in problem (L2), PDDF is by far the most effective as it is capable of converging with fewer sub-function evaluations, proving once more that when the function has a partially separable structure the proposed algorithm is particularly effective.
@@ 6 Conclusions
In this paper, we described and formally analyzed a penalty decomposition scheme for derivative-free optimization of coordinate partially separable finite-sum problems. The approach is proved to be well defined and possess global convergence guarantees under very mild standard assumptions. The results of the computational experiments indicate that the proposed penalty decomposition algorithm is capable of outperforming classical coordinate search methods and more elaborate mesh adaptive approaches in large scale structured problems. Moreover, the approach appears to be well-suited for parallelization; in scenarios where multi-processing is possible and the cost of computation is dominated by the evaluation of the black-box function, the proposed framework is shown to be superior also to a structure-informed smart implementation of the line-search based coordinate descent method. A deeper study in this setting might thus be of interest in future research.
The nice feedback obtained by our experiments opens up to interesting paths that could be explored in future research. First, working independently with each term in the objective should make the entire process less vulnerable to the presence of noise in a subset of the objectives. Moreover, the separability with respect to the sum terms is particularly nice in mixed setting where derivatives are available for some of the objectives [24]; indeed, gradient based optimization steps could be carried out w.r.t. blocks of variables associated with white-box terms; this would not be possible directly tackling the original formulation of the problem, where the same variables possibly appear in white and black box terms simultaneously. A further strength point of penalty decomposition approaches in general is that they are able to handle additional constraints with a limited overhead (see, e.g., the discussion in [19]): indeed, additional constraints can be moved to the penalty function and handled naturally within the sequential penalty scheme. Finally, the development of analogous penalty schemes based on the augmented Lagrangian function [5], instead of the simple quadratic penalty, could be considered; in particular, under convexity assumptions for the objective function, ADMM-type [28] methods might be appealing to solve problem (3).
@@
Matteo Lapucci: 0000-0002-2488-5486
Davide Pucci: 0000-0003-4424-9185
Marco Sciandrone: 0000-0002-5234-3174